OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). 7 -14 0 14

OpenStudy (anonymous):

\[slope=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

I'm really confused

OpenStudy (texaschic101):

this is the slope formula : slope = (y2 - y1) / (x2 - x1) now label your points (-4,8)....x1 = -4 and y1 = 8 (-3,-6)..x2 = -3 and y2 = -6 now we will sub...but be careful and watch your signs...thats the confusing part slope = (y2 - y1) / (x2 - x1) slope = (-6 - 8) / (-3 - (-4) slope = - 14 / (-3 + 4) slope = -14 / 1 or just -14 do you understand everything I did ? If not, please ask and I will explain the best I can